OpenStudy (anonymous):

Simplify: sin B + (cos2B/ Sin B)

OpenStudy (unklerhaukus):

is that a `cos^2(B)` or `cos(2B)`?

OpenStudy (anonymous):

Simplify: sin B + (cos^2 B/ Sin B)

OpenStudy (unklerhaukus):

factor out 1/sin B from both terms

OpenStudy (anonymous):

I don't understand how to

OpenStudy (unklerhaukus):

[ noting that the first term `sin B` is equal to `sin^2 B / sin B` ]

OpenStudy (unklerhaukus):

sin B + ( cos^2 B / sin B ) = sin^2 B / sin B + ( cos^2 B / sin B ) = 1 / sin B ( .... + .... ) = @Ambbiiee

OpenStudy (anonymous):

2 sin b/2 (cos b/2 + sin b/2 )=0

OpenStudy (unklerhaukus):

what? how

OpenStudy (anonymous):

Let me redo it

OpenStudy (unklerhaukus):

ok